When i try the solvers and choose elimination it automatically goes to substitution, i need both in order to complete the problem...i Dont know how so i cant complete the rest of the page!
3x - 5y = 7
2x + 5y = 13
Thank You!

Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

You add the two equations which gives 5x=20
x=4 now you do substitute 4 for x and solve for y
12-5y=7
subtract 12
-5y=-5
divide by -1
y=1

Answer by stanbon(75887)

3x - 5y = 7
2x + 5y = 13
----
Multiply thru 1st by 2.
Multiply thru 2nd by 3.
------
6x - 10y = 14
6x + 15y = 39
-----
Subtract and solve for "y":
25y = 25
y = 1
----
Solve for "x":
3x - 5y = 7
3x - 5*1 = 7
3x = 12
x = 4
---------
Cheers,
Stan H.
===========